### v2

# Enhancing Code LLM Training with Programmer Attention

Abstract.
Human attention provides valuable yet underexploited signals for code LLM training, offering a perspective beyond purely machine-driven attention. Despite the complexity and cost of collecting eye-tracking data, there has also been limited progress in systematically using these signals for code llm training. To address both issues, we propose a cohesive pipeline spanning augmentation and reward-based fine-tuning. Specifically, we introduce (1) an eye-tracking path augmentation method to expand programmer attention datasets, (2) a pattern abstraction step that refines raw fixations into learnable attention motifs, and (3) a reward-guided strategy for integrating these insights directly into a CodeT5 supervised fine-tuning process. Our experiments yield +7.16 in CodeBLEU on the CodeXGlue benchmark for code summarization, underscoring how uniting human and machine attention can boost code intelligence. We hope this work encourages broader exploration of human-centric methods in next-generation AI4SE.
Large Language Models, Eye Tracking, Code Comprehension
conference: ACM ACM International Conference on the Foundations of Software Engineering; 23-27 June, 2025; Trondheim, Norwaybooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwaycopyright: rightsretainedjournalyear: 2025conference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3728510isbn: 979-8-4007-1276-0/25/06ccs: Software and its engineering Automatic programmingccs: Information systems Document representation
1. Introduction
Programmers often exhibit selective attention when interpreting and modifying source code. Understanding these real-world attention patterns can deepen our grasp of code comprehension, improving tasks like summarization or completion (Hoq et al., 2024; Nijkamp et al., 2023). However, collecting such data (for example, via eye tracking or manual annotations) is expensive and time-consuming, limiting its large-scale use. Even when available, novel ways of integrating human attention signals into LLM training remain limited (Chen et al., 2021; Kou et al., 2024), so most AI-driven code tools rely on static artifacts alone. This gap underscores the need for data-augmentation strategies and specialized pipelines that systematically embed genuine programmer fixations into LLM-based code intelligence.
While LLMs like CodeT5 improve when fine-tuned on large code corpora, integrating human signals like eye-tracking attention patterns can enhance them further (Zhang et al., 2024; Kou et al., 2024). However, typical data augmentation often overlooks code’s cognitive demands or specific structures (Hamza et al., 2024; Chen et al., 2021). Consequently, even advanced LLMs may underutilize valuable programmer attention information (like scanpaths or reading orders), especially when such data is sparse. Effectively embedding these developer signals necessitates specialized training pipelines. To address this, we propose HumanLLM, a two-part approach primarily for code summarization (also tested on completion and translation). First, HumanLLM uses real eye-tracking data reflecting programmer reading to augment code. Second, it embeds these signals via reward-based fine-tuning of CodeT5 (Wang et al., 2021a), aligning the model with human fixations. This systematic use of human-centric patterns aims to bridge cognitive insights and automated code intelligence across tasks.
Although our attention patterns originate from code summarization studies and may not fully generalize to completion or translation, our experiments show that injecting these human-centric signals can notably improve model performance. For instance, we observed a up to +7.16 in CodeBLEU (Ren et al., 2020), +2.67 in Syntax, and +14.97 in Dataflow on the CodeXGlue dataset for code summarization, suggesting the potential of attention-driven methods to enhance software engineering workflows. This reinforces the vision that systematically integrating real programmer attention into code LLM training can unify cognitive insights with automated solutions across diverse AI4SE tasks.
2. Preliminaries
Our two-part code LLM fine-tuning pipeline leverages three key preliminaries to integrate developer attention:
Eye-Tracking for Programming.
Eye tracking captures fixations, where developers pause on tokens, and saccades, the rapid jumps between them, revealing real-time cognitive strategies in code reading (Minelli et al., 2015; Salvucci and Goldberg, 2000). These insights show how programmers parse syntax and identify relevant constructs, grounding our human-centered pipeline (Section 3).
Data Augmentation in AI for Code.
Traditional augmentation methods (e.g., token substitutions, AST edits) often overlook developers’ natural reading flows (Allamanis et al., 2018; Yu et al., 2022). By integrating actual eye-tracking data (Section 3.2), our approach captures these real patterns, helping models learn cognitively grounded cues that reflect authentic developer attention.
Code LLM Training.
Modern code LLMs typically follow Retrieval-Augmented Generation (RAG) or Supervised Fine-Tuning (SFT) (Chen et al., 2021; He et al., 2024). We employ a reward-guided SFT procedure, periodically injecting eye-tracking signals into the training loop (Section 3.3), so that model outputs better match actual programmer fixations.
3. Approach
Figure 1. Overview of our pipeline (HumanLLM): ① collects eye-tracking data (red) to capture real programmer attention; ② augments these fixations with AST-based adjacency (blue) and k-gram patterns; and ③ uses these human signals to guide a reward-based CodeT5 fine-tuning.
Figure 1 outlines our pipeline, which we refer to as HumanLLM, integrating eye-tracking data into a CodeT5 model in three stages. First, ① Data Collection gathers token-level fixation data from an open Java eye-tracking dataset (Karas et al., 2024) and supplements it with Java snippets from CodeXGlue [^1] https://github.com/microsoft/CodeXGLUE. Second, ② Augmentation enriches these fixations with AST-based semantic labels, adjacency expansions, k-gram patterns, and positional ordering (Sections 3.13.2). Finally, ③ Reward-Based Fine-Tuning periodically fuses these human-guided signals into CodeT5 training (Section 3.3), aligning the model with genuine programmer behavior.
3.1. Data Collection
We use the open eye-tracking dataset (Karas et al., 2024), which captures how 15 professional developers read and summarized 120 Java methods (30–50 lines each), logging token-level fixations and saccades. To enhance diversity, we incorporate Java samples from CodeXGlue. Each snippet provides code tokens (e.g., if, UserInput, 42), mapped to semantic tokens (Keyword, Identifier, Literal) via AST annotations. These fixation traces and mappings serve as the foundation for augmentation and reward-based training.
3.2. Human-Centric Augmentation of Code Tokens
Let $\mathbf{T}=\{t_{1},\dots,t_{n}\}$ be the tokens in a Java snippet, and let $\mathbf{F}\subset\mathbf{T}$ represent those actually fixated by developers. We construct an enriched set $\mathbf{F}^{\star}$ that encodes the following programmer attention information:
Semantic Labeling.
We parse each method’s AST and assign a label $L(t_{i})$ (e.g., Keyword, Literal) to each token $t_{i}\in\mathbf{T}$. This step ensures that the model recognizes syntactic roles rather than relying on raw text.
Adjacency-Based Expansion.
For each $t_{i}\in\mathbf{F}$, we include any tokens ±$\pm 3$ lines away that share $L(t_{i})$. Formally,
$\mathbf{F}^{\star}\;\leftarrow\;\mathbf{F}\;\cup\;\bigl{\{}\,t_{j}\,\big{|}\,|\text{line}(t_{j})-\text{line}(t_{i})|\leq 3\;\wedge\;L(t_{j})=L(t_{i})\bigr{\}}.$
This threshold of three lines stems from our analysis of the open eye-tracking dataset (Karas et al., 2024), showing that over 95% of attention shifts occur within this distance, indicating that developers often check nearby lines with similarly categorized tokens.
K-Gram Patterns.
Prior work on developer behavior (Karas et al., 2024; Zhang et al., 2024) shows that bigrams ($k=2$) and trigrams ($k=3$) reflect how programmers cluster code tokens. We detect these patterns in $\mathbf{F}^{\star}$, retain the top 20 for reward modeling, and assign each $p_{\ell}$ a numeric label. If a snippet matches both, we prioritize the trigram, breaking ties by frequency. Capturing these reading clusters models how developers group tokens during comprehension.
Positional Ordering.
Beyond adjacency expansions and k-gram patterns, we track each token’s reading index $\pi(t_{i})$ to align augmented tokens with actual scanpaths. This captures how developers encounter tokens in sequence, extending beyond simple locality. We record positions for up to 100 tokens per snippet, ensuring consistent fixation mapping. Each token $t_{i}\in\mathbf{F}^{\star}$ is labeled with $\bigl{(}p_{\ell}(t_{i}),\pi(t_{i})\bigr{)}$, encoding its semantic pattern and reading order for a more human-guided code representation.
3.3. Reward-Based Fine-Tuning with Programmer Attention
We integrate these augmented signals into a CodeT5 pipeline that periodically references the human-informed dataset $\mathbf{F}^{\star}$. After every $m$ mini-batches ($m=20$) of standard cross-entropy (CE) training, we sample a mini-batch from $\mathbf{F}^{\star}$, generate predictions $\hat{y}$, and compare them to the ground-truth labels $y^{\star}$. The resulting reward $\mathcal{R}(\hat{y},y^{\star})$ measures how well the model’s predicted sequences align with genuine developer fixations (e.g., matching k-grams or positional order). We then form a combined loss
$\mathcal{L}_{\mathrm{total}}\;=\;\mathcal{L}_{\mathrm{CE}}\;+\;\alpha\,\mathcal{R},$
where $\alpha$ regulates the strength of human guidance. By repeatedly backpropagating $\nabla\mathcal{L}_{\mathrm{total}}$, CodeT5 learns to minimize typical prediction error while conforming to real programmer attention patterns. Over multiple epochs, these reward passes steer the model to internalize adjacency relationships, k-gram fixations, and realistic scanpath ordering, thereby producing more human-aligned outputs (e.g., code summaries).
4. Experimental Design
We conduct two sets of experiments: (1) a basic Transformer analysis on an 80–20 split of our augmented eye-tracking dataset to assess whether the enriched signals are genuinely learnable, and (2) a CodeT5-based evaluation (using codet5-base [^2] https://github.com/salesforce/CodeT5) on three tasks in CodeXGlue for Java (summarization, completion, and Java-to-C# translation).
In the first setting, we train for a single epoch, tracking progress (0–100%) via the batch-size ratio to ensure the Transformer absorbs semantic (k-gram) and positional signals. In the second, we fine-tune CodeT5 with a reward-based loop (every 20 batches), adding an $\mathcal{R}$-driven term to cross-entropy that aligns predictions with real developer fixations. We fix seed 42 for data sampling and initialization, train on two NVIDIA A6000 GPUs (48 GB each), and use AdamW ($\text{lr}=5\times 10^{-5}$) with gradient checkpointing.
To balance diversity and runtime constraints, we sample 10% of each CodeXGlue dataset. We report BLEU, ROUGE-L, and METEOR for textual fidelity in code summarization, and CodeBLEU for deeper code-level accuracy (incorporating syntax and dataflow checks). This combination of metrics provides a comprehensive measure of both linguistic coherence and program correctness.
5. Experimental Results
We structure our analysis in three parts: (1) validating augmented signals across different adjacency windows and batch-size schedules, (2) showing how these signals improve CodeT5’s summarization (both NLP and code metrics), and (3) testing their transfer to completion and translation.
5.1. Learning with Human-Centric Augmentations
Adjacency Window Effects.
Figure 2 reports Precision, Recall, and F1 on a held-out test set for semantic and positional labels when expanding adjacency from 0 to 3 lines. In five of the six metrics, performance surpasses a baseline Transformer (dashed lines). Only Recall for positional labels lags slightly at the widest window, indicating that too much context can dilute the reading-order signal.
Figure 2. Impact of adjacency expansions (0 to 3 lines) on semantic (left) and positional (right) labels. Wider windows generally improve Precision, Recall, and F1, exceeding a baseline Transformer (dashed lines) in five of six metrics.
Batch-Size Ratio (Training Progress).
Figure 3 then examines how performance evolves during a single epoch, where the batch-size ratio on the horizontal axis reflects progress from 0% to 100% of that epoch. Larger ratios generally yield higher F1, Precision, and most Recall values, suggesting that the model benefits from encountering more human-centric samples earlier and more frequently.
Figure 3. Learning curves for semantic (left) and positional (right) labels over one epoch, shown via the batch-size ratio. Dashed lines represent the baseline Transformer. Larger ratios correlate with better test-set performance.
Together, Figures 2 and 3 confirm that adjacency expansions and sufficient training exposure help the Transformer internalize real programmer fixations.
5.2. Enhanced Code Summarization
We next assess whether these cues boost code summarization. Table 1 shows notable gains in BLEU, ROUGE-L, and METEOR, while Table 2 reveals parallel improvements in CodeBLEU, Syntax, and Dataflow (up by nearly 15 points). This suggests that real developer fixations help the model track variables and data dependencies, mirroring human comprehension.
Table 1. Summarization (NLP Metrics). HumanLLM integrates programmer attention into CodeT5, resulting in more fluent, context-aware summaries.
Pipeline BLEU ROUGE-L METEOR
Baseline 5.36 10.80 7.22
HumanLLM 10.32 14.91 12.88
Change +4.96 +4.11 +5.66
Table 2. Summarization (Code Metrics). By leveraging actual fixations, HumanLLM demonstrates improved structural understanding and variable handling.
Pipeline CodeBLEU Syntax Dataflow
Baseline 8.51 3.06 18.01
HumanLLM 15.67 5.73 32.98
Change +7.16 +2.67 +14.97
Figures 4 and 5 show that these improvements extend across different function types. HumanLLM outperforms the baseline in syntax and data flow for most categories, indicating the augmented signals help CodeT5 adapt to varied coding styles.
Figure 4. Overview of function types (left) and token-length distribution (right) for the ground-truth snippets in our dataset. The histogram on the left shows how frequently each function type appears, while the box plot on the right illustrates variability in token counts across different types.
Figure 5. Syntax (left) and Data Flow (right) scores by function type for Baseline vs. HumanLLM. HumanLLM generally outperforms the Baseline by better capturing control structures and variable interactions, aligning more closely with real developer fixations.
5.3. Task-Specificity for Completion and Translation
Table 3 shows results on code completion and Java-to-C# translation, where syntax alignment sees minor gains but CodeBLEU and dataflow do not improve substantially. This highlights the specialization of our method for summarization: these fixation patterns do not transfer easily to tasks that rely on different developer attention cues. Future extensions may require distinct eye-tracking protocols or reward definitions to optimize performance in completion, translation, or other AI4SE scenarios.
Table 3. Completion and Translation (Code Metrics). Summarization-centric attention signals offer limited gains for other tasks.
Task Pipeline CodeBLEU Syntax Dataflow
Completion Baseline 5.42 10.15 3.40
HumanLLM 4.77 10.46 1.45
Change -0.65 +0.31 -1.95
Translation Baseline 15.18 19.38 14.50
HumanLLM 14.94 18.96 14.38
Change -0.24 -0.42 -0.12
6. Future Work
Our observations using eye tracking data from professional developers show promising improvements in code summarization metrics. In future work, we will develop novel augmentation techniques to better leverage these signals and extend benefits to tasks such as code review, fault localization, and security auditing. By incorporating richer eye tracking data and flexible reward strategies, we aim to build AI driven coding assistants that capture developer focus and deepen code understanding.
7. Threat to Validity
Our study is based on an eye tracking dataset from professional Java developers, which may limit the generalizability of our findings to other languages and paradigms. The use of 10% of the CodeXGlue data may introduce sampling biases and, although we fix the random seed (42), some variation remains. Our reward scheme is designed for code summarization and may require adjustments for tasks such as bug detection, refactoring, or translation. Finally, the resource intensive nature of gathering eye tracking data calls for more efficient methods to leverage these signals.
8. Related Work
We briefly survey how code intelligence integrates human-centered signals, review neural approaches to code comprehension, and note why many LLMs overlook real programmer focus. Our method closes these gaps by infusing eye-tracking data into Code LLM training.
Human-Centered AI for Software Engineering.
Human input aids SE tasks (e.g., developer annotations (Hamza et al., 2024; Huang et al., 2023, 2025), or crowdsourcing (Lu et al., 2024)). Eye tracking reveals detailed reading strategies (Karas et al., 2024; Li et al., 2024b; Tang et al., 2024; Bansal et al., 2023a), yet most work applies these signals post hoc (Bansal et al., 2023c) or in narrow domains (Zhang et al., 2024). We directly integrate eye-tracking insights into LLM fine-tuning to capture genuine developer scanpaths.
Neural Code Analysis and Comprehension.
Deep learning powers code summarization, clone detection, and bug identification (Shi et al., 2022; Richter et al., 2022; Zhang et al., 2022a; Bansal et al., 2023b; Li et al., 2024a), often using tokens, ASTs, or graphs (Zhang et al., 2022b; Wang et al., 2021b; Pailoor et al., 2024; Acharya et al., 2025). These methods typically ignore real developer focus and rely on static artifacts. Our approach augments models with human attention, adding a more developer-centric lens to code comprehension.
LLMs for Code.
Models such as CodeBERT (Feng et al., 2020), Codex (Chen et al., 2021), and CodeT5 (Wang et al., 2021a, 2023) excel via large-scale pretraining and selective fine-tuning but often miss actual programmer fixations. Although some studies consider integrating human attention (Lai et al., 2020; Abulaish et al., 2024; Li et al., 2024b), efforts remain limited. Our reward-guided framework systematically aligns the model with genuine developer
9. Conclusion
We present a developer-attention–driven approach for fine-tuning Code LLMs, aligning model outputs with real programmer fixations using eye-tracking data. By augmenting these signals with semantic labels and positional cues and applying reward-based fine-tuning to CodeT5, we achieve notable gains in textual and code-specific metrics. Our results underscore the potential of integrating human insight with AI, fostering deeper collaboration in AI4SE.